### v1

# Vulnerabilities of quantum key distribution systems in visible range

Abstract
Spectral loopholes in the 1000-2100 nm range have been the focus of attention of quantum hacking research aimed at searching and elimination of vulnerabilities in quantum key distribution systems. In this work, we concentrate on shorter wavelengths ranged from 400 nm up to 800 nm and experimentally study spectra of insertion losses for a number of fiber optical components to assess potentials for their implementation as countermeasures. We show that efficiency of the elements commonly used as countermeasures against hacking attacks in the telecom range can be significantly impaired in the visible range. It is found that the vulnerabilities detected in this range play an increasingly important role for hacking strategies such as the induced-photorefraction attack whose efficiency improves when the wavelength decreases.
preprint: APS/123-QED
I Introduction
The security of quantum key distribution (QKD) systems hinges upon the fundamental laws of physics. Over the past decades, significant progress has been made in theoretical security proofs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. However, many of these proofs rest on the assumption of perfect and trusted devices. Such trust assumption imposes a rather strong security requirement which is generally difficult to implement in practice. In real situations imperfect devices in QKD systems would introduce vulnerabilities that can be exploited by eavesdroppers to extract additional information about the transmitted bits during their attacks as it is considered in [11, 12, 13, 14, 15, 16]. Device-independent (DI) [17, 18, 19, 20, 21] and semi-device-independent [22, 23, 24, 25] protocols have emerged as a prominent area of research that provides a way to prevent side-channel attacks [26]. On the opposite side, another rapidly developing field which is often referred to as quantum hacking has recieved a considerable amount of attention [27].
Emerging approaches in quantum hacking pose novel challenges for security analysis of QKD systems. Numerous vulnerabilities have been identified in prior works [28, 29, 30], underscoring the importance of enhancing security measures. Eavesdropper strategies in quantum hacking can broadly be classified under two basic types, active and passive, as well as combinations thereof. Active strategies involve the manipulation of optical parameters within the hardware, whereas passive strategies solely facilitate the acquisition of additional information.
A quintessential illustration of a passive attack is the Trojan-horse attack which is also known as the large pulse attack [31, 32]. In this scenario, the eavesdropper (Eve) employs scanning pulses introduced into the systems of legitimate users (Alice and Bob) via the quantum channel, allowing for the extraction of distributed sequences from the reflected portions of the eavesdropper pulses back into the quantum channel. The key factors that determine the efficiency of this attack involve the transmittance of elements in two directions, the distinguishability of states along with countermeasures utilized in the system. An important point is that all these factors require the eavesdropper to define the optimal spectral range of wavelengths employed to perform attack.
The well-known detector blinding attack (DBA) [33, 34, 35] exemplifies an active attack. The idea underlying the attack is that, under the action of high-power laser radiation (as compared to typical intensities used in QKD), the operating mode of single photon detectors (SPDs) based on avalanche photodiodes can be changed from the Geiger mode into the linear operating mode, so that the eavesdropper can manipulate the clicks registered by the SPDs. The efficiency of this DBA based quantum hacking approach crucially depends on the spectral sensitivity of the avalanche photodiode, the transmittance of optical elements connecting the quantum channel with the single photon avalanche diode (SPAD), and the countermeasures implemented into the QKD system. So, it turned out that, similar to the Trojan-horse attack, the spectral range is one of the factors that govern success of the DBA attack.
Previous studies [36, 37, 38, 39] motivated by the THA, have mainly been focused on transmittance of fiber optical elements utilized in QKD systems at wavelengths ranged between 1000 nm and 2100 nm. For the 600-1000 nm range, the experimental results were reported in Refs. [39, 40].
But, for the above discussed DBA and the laser damage attacks (LDA) that exploit temperature dependence of the transmittance of fiber-optical elements [41, 42, 43], shorter wavelengths may also be used. At shorter wavelengths, low insertion losses of elements employed for countermeasures may lead to unnoticed and thus potentially successful delivery of radiation injected by Eve to the elements under attack.
New vulnerabilities at different wavelengths have been the subject of intense recent studies [44]. The induced-photorefraction attack (IPA) [27, 45] represent attacks that will do better in the visible spectral range [46]. In addition, this attack is claimed to be applicable for hacking measurement device independent (MDI) QKD systems [27].
The main targets of the IPA are phase and amplitude modulators based on photo-refractive crystals, where variations in refractive index or transmittance can be induced by light. The IPA strategy enables Eve to manipulate the characteristics of quantum states, i.e. mean photon number, at the sender’s side using laser radiation. The efficiency of these attacks improves as the radiation wavelength becomes shorter [46]. So, it is important to perform spectral measurements of fiber components used in QKD systems at shorter wavelengths so as to identify and mitigate vulnerabilities exposed by the attack. In particular, the 400-800 nm spectral range has not been received a proper attention yet. In this paper, our goal is to fill the gap.
Note that, according to Ref. [46], successful implementation of the IPA requires low intensities starting from 3 nW. A comparison between these intensities and the maximal intensity permissible for insertion into a fiber without causing damage, which is typically around 9 W (this value is applicable to quartz single mode fibers without power damage level doping commonly used in fiber communications and may vary among manufacturers) leads to the conclusion that attenuation needed for passive defence against IPA for single pass through optical elements can roughly be estimated at about 100 dB. Thus, in order to assess the feasibility of the attack, it is necessary to measure and analyze insertion losses of individual optical elements on the side of Alice.
To this end, we study the insertion losses of different optical elements in the 400–800 nm spectral range, demonstrate the existence of potential vulnerabilities and suggest potential countermeasures. The paper is organized as follows.
In Sec. II we describe our experimental technique employed to measure spectra of insertion losses of optical elements at wavelengths ranged from 400 nm to 800 nm. The experimental data for insertion losses measured for a variety of optical attenuators, wave-division multiplexing (WDM) and isolating components are presented in Sec. III. Finally, in Sec. IV, we discuss and summarize our results.
II Experimental setup and method
Figure 1 shows the optical scheme of our experimental setup used to measure the insertion losses of optical elements in the visible range. In this scheme, a broadband plasma light source (XWS65, ISTEQ) indicated as the broadband light source (BLS) is utilized to produce light with wavelengths ranged from 250 nm up to 2500 nm. Referring to Fig. 1, the radiation from this source is directed into a single-mode fiber (OF) passing through a set of lenses (L), filters (F) and the objective (O) which is positioned on a three-axis stage to meet coupling and spectrum requirements. After the fiber, the focused light propagates through either a fiber or a fiber with the optical elements under test and is measured by the spectrometer (USB4000-UV-VIS-ES, Ocean Optics) with the 1.5 nm resolution.
Figure 1: Optical scheme for insertion losses measurements in the visible range. BLS is broadband light source, L is lens, F is set of filters, O is objective, OF is single mode optical fiber, EUT is element under test, S is spectrometer
Figure 2: Insertion losses measured in the 400-800 nm range for (a) electro-mechanical and (b) electro-optical variable optical attenuators (blue and red solid lines indicate the zero-voltage curve and the case where the applied voltage is 5 V, respectively); (d) absorption based fixed attenuators (the curves for the first 20 dB FOA, second 20 dB FOA and 10 dB FOA are indicated by black, red and blue solid lines, respectively); (c) scattering based fixed attenuators principles (the curves for 20 dB FOA, 15 dB FOA and 5 dB FOA are indicated by black, red and blue solid lines, respectively).
An important point is that the spectrometer is equipped with its own multimode fiber with an SMA905 connector. So, in order to align the single-mode and multimode fibers, we have used the FC-SMA905 connector.
Due to variations in efficiency of coupling to the single-mode fiber, losses due to absorption by the filters and technical restrictions imposed by the spectrometer, the measured spectrum of the broadband light source (BLS) was narrowed to the 400-800 nm range. Insertion losses of the optical element under test was calculated as follows:
$\alpha_{\mathrm{dB}}(\lambda)=-10\log_{10}\bigl{[}P_{\mathrm{mes}}(\lambda)/(P_{\mathrm{ref}}(\lambda)T_{\mathrm{f}}(\lambda))\bigr{]},$ (1)
[…]
III Experimental results
III.1 Optical attenuators
In this section, we present the results for two variable optical attenuators (VOAs) that differ in the mode of operation along with six different fixed optical attenuators (FOAs).
In QKD systems optical attenuators (OA) are typically employed to reduce the intensity of laser pulses to the single photon level. Additionally, OAs are of importance in reducing the probability of wavelength-dependent attacks successful realization. Consequently, there is a requirement to avoid pronounced variations of spectral insertion loss properties of OAs across a broad wavelengths range.
In QKD systems, the insertion losses of VOAs are typically adjusted by applying voltage. This implies that during the key distribution process, the attenuation value can be tuned in the range between the maximum and minimum insertion losses. To address this variability, we have evaluated the maximal and minimal insertion losses settings for one electro-mechanical (EM) and one electro-optical (EO) model of VOA. According to the manufacturer datasheets, at the wavelength $\lambda=1550$ nm, the minimal values of insertion losses without power (the field free regime without applied voltage) are 0.77 dB and 0.80 dB for EO VOA and EM VOA, respectively. When the applied voltage is 5 V, the corresponding measured values of insertion losses for EO VOA and EM VOA are 36.0 dB and 43.4 dB.
Figure 3: Insertion losses measured in the 400-800 nm range for (a) CWDM1; (b) CWDM2; (c) CWDM3 and (d) DWDM (black, blue, and red solid lines indicate insertion losses measured for light propagating along three different directions: from the ref-port to the 1550-port, from the com-port to the ref-port and from the com-port to the 1550-port, respectively)
Figures 2(a) and (b) present wavelength dependencies of insertion losses measured in the 400-800 nm range for EM VOA and EO VOA at two values of applied voltage: 0 V and 5 V.
In the zero-voltage case, it turned out that the measured insertion losses are higher than the values reported in the datasheet. For EM VOA and EO VOA, the mimimal losses 10.7 dB and 21.8 dB are reached at the wavelengths 761 nm and 557 nm, respectively.
Referring to Figs. 2(a) and (b), when the applied voltage is 5 V (blue lines), the losses are shown to vary in the range between 13.7 dB and 30 dB for EM VOA, whereas, for EO VOA, they are ranged from 23.7 dB to 50 dB. Note that these maximum values of losses are approximate as they correspond to the noise level of our experimental setup and depend on the number of neutral filters.
Insertion losses measured in relation to the wavelength for six FOAs are shown in Figs. 2(c) and (d). For three absorption based FOAs (see Figs. 2(c)), according to the values reported by the manufacurers, the insertion losses of one FOA is 10 dB, whereas the losses for the other two FOAs are the same and equal to 20 dB. For the scattering based FOAs (see Figs. 2(d)), there are three different reported values of losses at the operating wavelength ($\lambda=1550$ nm): 5 dB, 15 dB and 20 dB. Our experimental results are in good agreement with these values.
The insertion losses of the absorption based FOAs measured in the 400-800 nm range are shown in Fig. 2(c). It can be seen that, for each FOA, there are two transmittance windows where the insertion losses are noticeably lower than the value at the operating wavelength indicated in the legend. The window ranged between 400 nm and 430 nm is located in the vicinity of local minima for the two 20 dB attenuators with $\alpha_{\mathrm{dB}}\approx 8.1$ and $\alpha_{\mathrm{dB}}\approx 8.6$ at the wavelength $405$ nm (for the 10 dB attenuator, the corresponding value is $\alpha_{\mathrm{dB}}\approx 2.7$) Similarly, the other window in the 750-800 nm wavelengths range is close to the local minima at $\lambda\approx 780$ nm with the inserion losses estimated at $\alpha_{\mathrm{dB}}\approx 3.4$ and $\alpha_{\mathrm{dB}}\approx 4.3$ and $\alpha_{\mathrm{dB}}\approx 0.9$.
Figure 2(d) shows the wavelength dependencies of the insertion losses measured for the scattering based FOAs. It is seen that, for the 20 dB attenuator, the insertion losses insignificantly change with wavelength varying from 27 dB to 25 dB. Similar conclusion applies to the 15 dB (5 dB) attenuator, where variations are between 15.6 dB (11.2 dB) and 13 dB (10.8 dB).
III.2 Wavelength-division multiplexing components
In QKD systems, wavelength-division multiplexing (WDM) components can be used as countermeasures by suppressing radiation with wavelengths outside the range utilized by the communication channel between Alice and Bob. In the simplest case, the key elements of the internal design of fiber WDM component are thin films deposited on a substrate. Ideally, at the working wavelength, light will enter the common port (the com-port) and, after passing through the thin films and the substrate, will propagate to the corresponding port of the operating wavelengths grid. Other wavelengths are directed to the reflecting port (the ref-port). This port might be connected with a watchdog detector that monitor Eve’s probing pulses. The latter assumes that the sensitivity range of the watchdog detector is broad enough to cover wavelengths of potential attacks.
In practice, there is a number of factors limiting the efficiency of the WDM-components. These include the substrate transmittance as well as the number, ordering and the thicknesses of the thin films. Differences between the perfectly designed mode of operation and its practical implementation may lead to side channels available to the eavesdropper. In our previous study [37], the WDM-components are found to have the transmittance windows outside the telecom range in agreement with the theoretical calculations reported in [47].
Similar to Ref. [37], we consider both coarse WDM (CWDM) and dense WDM (DWDM) components. Specifically, insertion losses in the 400-800 nm range are measured for the four filters: three CWDM-components from the same manufacturer where two of them belong to the same production batch (CWDM1 and CWDM2) and one DWDM-component. The results are shown in Fig. 3.
Figure 4: Insertion losses measured for (a) the dual stage isolator with light propagating from the output port to the input port; (b) the circulator (blue (red) solid lines represent insertion losses measured for light propagating from the third (fourth) port to the first (second) port).
In addition to the com-port and the ref-port, the WDM components have the 1550-port corresponding to the operating wavelength $\lambda=1550$ nm. Referring to Figure 3, when the probing light propagates from the com-port to the 1550-port, each WDM-component reveals the ranges of wavelengths where the insertion losses are within the interval between $10$ dB and $20$ dB. Minimal values of insertion losses are presented in Table 1. Note that, for the WDM-components under consideration, the losses in the telecom range at non-working wavelengths are about $30$ dB.
Component
$\alpha_{\mathrm{min}}$ (dB)
$\lambda_{\mathrm{min}}$ (nm)
CWDM1
10.0
680
CWDM2
10.7
679
CWDM3
9.7
662
DWDM
10.3
665
Table 1: Minimal insertion losses and corresponding wavelengths for WDM components.
From Fig. 3, it can be seen that, for the light propagating from the com-port to the ref-port, the insertion losses vary between 10 dB and 30 dB which is close to the values for the above discussed direction from the com-port to the 1550-port. For the case of light propagation from the 1550-port to the ref-port, it turned out that the losses are close to the noise level. Note that, when directions of light propagation are reversed, changes in insertion losses appear to be well below accuracy of measurements (the results of these measurements are omitted) and thus nonreciprocity effects have not been observed in our experiments.
III.3 Isolating components
Isolating elements are typically utilized to reduce reflections back into laser resonators. Similar to the WDM components, in QKD systems, these elements may serve as countermeasures against attacks.
In this section, we examine two elements: the dual stage isolator and the circulator. Since we are mainly interested in countermeasures against the IPA attack, the case of backward propagation of injected light where probing pulses travel from the output port to the input port will be our primary concern.
For the isolator, the spectrum of insertion losses is shown in Fig. 4(a). At the working (operating) wavelength (1550 nm), the measured insertion losses for light propagating in backward direction is about 65 dB.
From the curve depicted in Fig. 4(a), the values of measured insertion losses are mostly below the noise level that can be detected in our experimental scheme (the dynamic range of our measurements is about $50$ dB). But, in the 772-800 nm range with the minimal losses 44.5 dB reached at $\lambda=784$ nm, the insertion losses appear to be lower than the above value at the operating wavelength.
Similar to isolators, in circulators, losses for light propagating from one port to another (forward direction) are low, whereas the light is nearly blocked when the direction of propagation is reversed (backward propagation). In contrast to the isolators, the circulators have several ports (in our case, the circulator has four ports) that opens up additional possibilities for applications in more complicated schemes. For a number of different reasons, in real devices, the insertion losses for light propagating in the backward direction will generally be wavelength dependent and may deviate from the value reported by manufacturers.
According to the datasheet for the circulator, at $\lambda=1550$ nm, the insertion loss for backward direction is about 50 dB.
Insertion losses spectra measured for the cases where light propagates from the third (fourth) port to the first (second) bypassing the second (third) port are presented in Fig. 4(b). It can be seen that in the 400-506 nm range the losses are lower than the reported value at the working wavelength. Minimal value of insertion losses is 31.1 dB at 419 nm.
Owing to limitations of our experimental setup, we have not detected any noticeable transmission of light for backward propagation between adjacent (neighbouring) ports (for instance, propagation from the second port to the first port). In this case, the results are well below the noise level and bear close resemblance to the curves measured for the isolator.
IV Discussion and conclusion
In this paper, we have studied spectra of insertion losses of various fiber optical elements such as optical attenuators (see Fig. 2), WDM and isolating components (see Figs. 3 and 4, respectively) in the visible light range with wavelengths from $400$ nm to $800$ nm. Both the optical elements and the spectral range are important for understanding vulnerabilities of real-world QKD systems to hacking attacks whose efficiency being wavelength dependent may noticeably improve in the visible range. Our study is motivated by one of such attacks known as the induced-photorefraction attack [27, 45, 46].
The optical scheme behind our experimental technique enables us to perform measurements of the insertion losses with approximately $50$ dB dynamic range. It is simple to use and low-cost when compared with analogous schemes discussed, e.g., in Ref. [30]. An important point is that, by using suitably modified methods presented in Ref. [38], the proposed scheme can be improved so as to extend both the wavelength range and the dynamic range of measurements. As far as certification of QKD systems is concerned, this is a feasible and easy-to-use scheme that can be employed for characterization of optical elements insertion losses.
Our experimental data show that, OAs implemented into QKD systems as countermeasures against attacks that utilize probing pulses in the visible range require careful considerations. For VOAs, it turned out that, at certain non-vanishing voltages, the insertion losses in the visible range are lower than the losses at the operational wavelength (see Figs. 2(a)-(b)). The results for the absorption based FOAs (see Fig. 2(c)) indicate the transmittance windows that can be exploited by the IPA attack. By contrast, the insertion losses of the scattering based FOAs (see Fig. 2(d)) do not reveal noticeable changes with wavelength. So, this type of OAs is preferable as additional countermeasure against the IPA attack.
The measured spectra of all the WDM-components (see Fig. 3) have transmittance windows where insertion losses are not high enough to prevent the IPA. It is also shown that the insertion losses for light propagating from the com-port to the ref-port are ranged between 10 dB and 30 dB. Since this direction is used for active monitoring purposes, such losses combined with low efficiency of InGaAs photodiodes in the visible range will make the monitoring ineffective. One way around this problem is to use Si-based photodiodes as the watchdog detectors against optical probing pulses in the 400-800 nm range.
From the insertion loss spectrum shown in Fig. 4(a), the dual stage isolator delivers at least 44 dB insertion losses. Presumably, such losses can be attributed to absorption of magneto-optical material used in the isolator. Similar insertion losses were measured in the circulator for the case of backward propagation of light between adjacent ports.
Insertion losses for the light passing from the third port to the first port plotted against wavelength in Fig. 4(b) show that the well-known broadband filtering scheme [48] based on a Bragg grating connected to the second port of the circulator becomes ineffective in the 400-506 nm range. Thus it cannot be employed as a reliable tool to counter the IPA.
We summarize with the remark that the scattering based FOA characterized by nearly wavelength independent losses and utilized in combination with a couple of the isolating components will deliver at least 100 dB insertion losses which is enough to counter successful realization of the IPA.
Acknowledgements.
The study is partially funded by the Ministry of Education and Science of the Russian Federation (Passport No. 2019-0903).